OpenStudy (anonymous):

Find the vertices and foci of the hyperbola with equation x^2/4 - y^2/60 = 1 1) Vertices: (± 2, 0); Foci: (± 8, 0) 2) Vertices: (± 8, 0); Foci: (± 2, 0) 3) Vertices: (0, ± 8); Foci: (0, ± 2) 4) Vertices: (0, ± 2); Foci: (0, ±8)

OpenStudy (anonymous):

for your reference: standard form of hyperbola (horizontal axis) (x-h)^2 / a^2 - (y-k)^2 / b^2 = 1 asymptote slope m = ± b/a standard form of hyperbola (vertical axis) (y-k)^2 / a^2 - (x-h)^2 / b^2 = 1 asymptote slope m = ± a/b (h,k) => center a: distance between center and vertices c^2 = a^2 + b^2 c: distance between center and focus .........................................

OpenStudy (anonymous):

Find an equation in standard form for the hyperbola with vertices at (0, ±2) and foci at (0, ±7). y^2/4-x^2/45=1

OpenStudy (anonymous):

Did that help ?

OpenStudy (anonymous):

it did but i still don't know what i'm doing

OpenStudy (anonymous):

what would be the center h,k

OpenStudy (anonymous):

im not really sure thats a good question lol

OpenStudy (anonymous):

yeah, i'm having trouble with two more problems like this

OpenStudy (anonymous):

i can help a little

OpenStudy (anonymous):

thats all i need

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what would be the first step to solving this